OpenStudy (jkbo):

The midpoint of uv is (5, –11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V.

OpenStudy (anonymous):

Work backwards from the midpoint formula. M(5, -11) = ((5 + x_2)/2, (-11 + y_2)/2)

OpenStudy (lalaly):

for the points (x1,y1) and (x2,y2) the coordinates of the midpoint are\[(\frac{x1+x2}{2}, \frac{y1+y2}{2})\]so u in ur question the midpoint is (5,-11) \[(5,-11)=(\frac{3+x2}{2}, \frac{5+y2}{2})\] now for the x coordinate of V\[5=\frac{3+x2}{2}\]solve for x2 and for the y coordiante of V\[-11=\frac{5+y2}{2}\]solve for y2

OpenStudy (jkbo):

thanks